### v1

# Velocitune: A Velocity-based Dynamic Domain Reweighting Method for Continual Pre-training

Abstract
It is well-known that a diverse corpus is critical for training large language models, which are typically constructed from a mixture of various domains. In general, previous efforts resort to sampling training data from different domains with static proportions, as well as adjusting data proportions during training. However, few methods have addressed the complexities of domain-adaptive continual pre-training. To fill this gap, we propose Velocitune, a novel framework dynamically assesses learning velocity and adjusts data proportions accordingly, favoring slower-learning domains while shunning faster-learning ones, which is guided by a scaling law to indicate the desired learning goal for each domain with less associated cost. To evaluate the effectiveness of Velocitune, we conduct experiments in a reasoning-focused dataset with CodeLlama, as well as in a corpus specialised for system command generation with Llama3 and Mistral. Velocitune achieves performance gains in both math and code reasoning tasks and command-line generation benchmarks. Further analysis reveals that key factors driving Velocitune’s effectiveness include target loss prediction and data ordering.
Velocitune: A Velocity-based Dynamic Domain Reweighting Method for Continual Pre-training
Zheheng Luo♠†thanks: The first two authors contribute equally. Work done during an internship at Microsoft. Correspondence: zhehengluo@gmail.com;  xinzhang3@microsoft.com     Xin Zhang11footnotemark: 1     Xiao Liu Haoling Li♢†     Yeyun Gong     Chen Qi     Peng Cheng The University of Manchester Microsoft Tsinghua University
1 Introduction
Datasets used for pre-training language models (LMs) are typically composed of texts of various meta-attributes like source and focus, referred to as different domains (Du et al., 2022; Azerbayev et al., 2023; Computer, 2023). The distinct characteristics of data from these varying domains, such as focus, quality and quantity, affect the downstream performance of LMs differently (Rozière et al., 2024; Li et al., 2024). Consequently, numerous studies have explored the optimal combination of data from multiple domains to enhance LM performance.
Figure 1: The overall pipeline of Velocitune. Initially, a proxy model is trained using the original domain weights on a subset of the data. Following this, the initial loss is collected by evaluating the base model, while the target loss is determined by extrapolating the evaluation loss of the proxy model. In the second phase, we calculate the learning velocity by rescaling the learning progress between the initial and target losses. This learning velocity is then used to update the domain weights effectively.
Llama3 (AI@Meta, 2024), GLaM (Du et al., 2022), and Lemma (Azerbayev et al., 2023) employ heuristic methods to iteratively test different ratios by training multiple proxy models and selecting the mixture that demonstrates the best downstream performance. However, these heuristic approaches demand costly large-scale experiments for effective exploration. As a result, recent research focuses on learning an optimal ratio by dynamically adjusting weights during proxy model training (Xie et al., 2024; Fan et al., 2023). For example, Doremi (Xie et al., 2024) implements a method in which a small reference model is initially trained, followed by training a small proxy model using group distributionally robust optimization (Group DRO) (Sagawa et al., 2019) to obtain domain weights.
Domain-adaptive continual pre-training, while sharing some similarities with from-scratch training, presents unique challenges that limit the effectiveness of existing domain reweighting methods. Many existing methods utilise small proxy models to estimate optimal domain weights, which are subsequently transferred to a larger model (Xie et al., 2024; Fan et al., 2023; Azerbayev et al., 2023). However, this approach poses challenges in domain-adaptive continual pre-training, as smaller versions of the base model often do not exist, making it difficult to estimate weights from these proxy models. Another challenge is how to leverage the learning status. Previous methods rely on the distance between the model’s current loss and a target domain loss (Xie et al., 2024; Xia et al., 2023). Nevertheless, this distance-based approach can result in uneven domain emphasis, as domains with larger loss disparities may disproportionately influence learning. This can exacerbate imbalances across domains.
To address these issues, we introduce a novel framework, Velocitune, centered on the concept of learning velocity. In contrast to previous approaches that leverage the distance between the current loss and a target loss, Velocitune more accurately capture how effectively models learn across each domain by establishing learning velocity. During training, domains exhibiting slower learning velocities are given increased weights, while those with faster velocities receive reduced weights, promoting equilibrium in learning progress. To quantify learning velocity, it is crucial to determine both the initial loss, reflecting the model’s already learnt expertise, and the target losses, indicating the desired learning goal for each domain. For a more cost-effective estimation of target losses, we adopt the Chinchilla scaling law (Hoffmann et al., 2022). Instead of using a small model, we predict target reference losses by leveraging models trained on sub-sampled data.
We evaluate the performance of Velocitune in two settings: continual pre-training CodeLlama 7B (Rozière et al., 2024) on a math and coding reasoning dataset as well as Llama3 and Mistral (Jiang et al., 2023) on a system command knowledge corpus. Velocitune demonstrates an average improvement of 1.6% across seven math tasks and 3.8% across two coding tasks compared to baseline models trained with default weights. In addition, Velocitune outperforms the baselines on Llama3, showing improvements of 5.3% and 3.1%, and on Mistral, with gains of 4.4% and 2.6% on the CmdGen-Nvidia and CmdGen-AMD tasks, respectively. Further more, we conduct an in-depth ablation study to identify key factors contributing to the observed improvements.Our findings indicate that, beyond the contribution of reweighted data ratios, the sequence of data ordering might play a significant role in the effectiveness of Velocitune. The results also show that the predicted target loss are critical for the success of the Velocitune. The contributions of this work can be summarised as follows:
Introducing Velocitune, a novel framework designed to adjust data ratios in continual pre-training. It establishes learning velocity to more accurately assess the extent of models’ learning in domains and leverages a scaling law to minimise the costs.
Demonstrating through extensive experiments that Velocitune enhances downstream performance in two continual pre-training settings.
Providing a detailed analysis revealing that data ordering, reweighted data ratios, and predicted target loss significantly contribute to Velocitune’s effectiveness.
2 Domain-adaptive continual pre-training with Velocitune
In this section, we present a detailed explanation of Velocitune. Our approach focuses on adjusting domain weights based on learning velocity. To quantify it, it is essential to determine both the initial and target losses for each domain. To achieve a cost-effective estimation of target losses, we adopt the Chinchilla scaling law to predict it using models trained using sub-sampled data. During training, we periodically assess learning velocity, increasing the sampling weights of slower domains while reducing weights of faster ones. This adjustment ensures more balanced learning across domains. The methodology is detailed in § 2.1.
2.1 The Velocitune algorithm
Setup
Consider training a language model (LM) on a dataset $S$ consisting of $n$ distinct domains, denoted as $D^{1},D^{2},\dots,D^{n}$. Let ${\mathbb{S}}={S^{1},S^{2},\dots,S^{n}}$ represent the subsets of data corresponding to each domain, where each $S^{i}$ contains both training and evaluation data, denoted as $S_{\text{train},i}$ and $S_{\text{eval},i}$, respectively. The domain weight $w\in\Delta_{n}$ represents the sampling weight assigned to domains.
Target Estimation
We first compute the initial loss ${\bm{L}}_{\text{init}}$. This involves evaluating the LM on the full evaluation set ${\mathbb{S}}_{\text{eval}}$, providing an accurate measure of the model’s learnt expertise in each domain. The target loss for each domain is derived using the scaling law (Kaplan et al., 2020). Specifically, in the absence of models with varying sizes, we apply the Chinchilla scaling law by fisrt defining the total dataset size and then training the model on a subsample using the default domain weights, which correspond to the ratio of tokens in each domain. Throughout the training process, multiple checkpoints are saved, and the evaluation losses from these checkpoints are used to fit the scaling law parameters. Finally, we take the fitted function to predict the loss the model could reach for using the entire dataset. More detailed analysis of prediction errors is provided in Appendix B.
Velocity-Guided Training
After obtaining the initial losses and target losses in each domain, our objective is to iteratively update domain weights, denoted by ${\bm{w}}_{t}$, to minimize the re-weighted training loss effectively. The optimization problem is formulated as follows, encompassing two sequential steps:
(1). Minimise the weighted training loss:
$\min_{\theta}\sum_{t=1}^{T}\sum_{i=1}^{k}w_{t}[i]\cdot\ell_{t}^{i}(\theta)$
where $\ell_{t}^{i}(\theta)$ denotes the training loss at step $t$ for domain $D_{i}$ with model parameters $\theta$. The $w_{t}[i]$ refers to the normalized weight assigned to domain $D_{i}$ at step $t$.
(2). Minimise the weighted sum of learning velocities:
$\min_{w\in\Delta^{k}}\sum_{i=1}^{k}w_{t}[i]\cdot V_{t}[i],$
The first step is to minimise the language modeling Negative Log Likelihood(NLL) loss given the domain weights. The second step focuses on dynamic adjustment such that the weight in the slowest learning velocity domain is maximised. By ensuring a focus on the most challenging scenarios, this helps in improving overall balance during training. Unlike the Group DRO optimization in Doremi (Xie et al., 2024), which updates domain weights based on the loss difference of each data point between a proxy model and a trained reference model, our method does not rely on a fully trained reference model. Consequently, the loss of each data point on the trained model is unobtainable. Instead, after a set number of training steps, we update the domain weights by first calculating the learning velocity, which reflects how quickly the model is learning in a given domain. The learning velocity is defined as:
$V_{t}[i]=\frac{\ell_{t}^{i}(\theta,S^{\prime}_{\text{eval},i})-\ell_{\text{target}}^{i}}{\ell_{\text{init}}^{i}-\ell_{\text{target}}^{i}}$
Then the domain weights are updated every $m$ steps as:
$w_{t}\leftarrow\frac{w_{t-m}[i]\cdot\exp(V_{t}[i])}{\sum_{i=1}^{k}w_{t-m}[i]\cdot\exp(V_{t}[i])}$
This formulation dynamically adjusts domain weights based on model’s learning velocities in domains, ensuring an equitable comparison of learning efficiency across domains with varying initial and target losses. By doing so, we prioritize learning in domains where progress towards the target loss is most promising, thereby enhancing overall model performance. In summary, this approach dynamically rebalances the learning process across multiple domains, adjusting weights in response to observed rate of learning progress, and steering them towards achieving optimal and uniform performance improvements. The complete Velocitune algorithm is summarized in Algorithm 1.
Require: Training data of $k$ domains $D_{1},D_{2},\cdots,D_{k}$, validation data $D^{\mathrm{val}}_{1},D^{\mathrm{val}}_{2},\cdots,D^{\mathrm{val}}_{k}$, initial data loading weights $\textbf{w}_{0}\in\mathbb{R}^{k}$ initialise as uniform distribution, upper bound loss $\ell_{\mathrm{U}}\in\mathbb{R}^{k}$, lower bound loss $\ell_{\mathrm{L}}\in\mathbb{R}^{k}$, LM loss $\mathcal{L_{T}}$ at training time $T$, evaluation per $m$ steps, model parameters $\theta$
for $t=1,\cdots,T$ do
       if $t\mod m=0$ then
             $\Delta_{t}[i]\leftarrow\textnormal{{Clamp}}\left\{0,\frac{(\ell_{t}^{i}(\theta,S^{\prime}_{\text{eval},i})-\ell_{\mathrm{target}}[i])}{(\ell_{\mathrm{init}}[i]-\ell_{\mathrm{target}}[i])},1\right\}$triangleright Rescale training progress
             $w_{t}\leftarrow$ UpdateWeight($w_{t-m}$, $\Delta_{t}$) $\triangleright$ Update data loading proportion
       end if
      Sample a batch of data $\mathcal{B}$ from $D_{1},D_{2},\cdots,D_{k}$ with proportion $w_{t}$;
       Update $\theta$ with $\mathcal{L}(\theta,\mathcal{B})$
end for
Function UpdateWeight($w$, $\Delta$)
       $\alpha\leftarrow w\cdot\exp\left(\Delta\right)$triangleright Calculate the unnormalized weights
       $w\leftarrow\frac{\alpha}{\sum_{i}\alpha[i]}$ return $w$triangleright Renormalize and smoooth the data loading proportion
Algorithm 1 Velocitune
3 Experiment
In this section, we apply Velocitune in continual pre-training CodeLlama-7B on the reasoning dataset and Llama3-8B as well as Mistral-7B on the system knowledge dataset.
Figure 2: Domain weights dynamic of Velocitune in training CodeLlama 7B on Reasoning.
Model Math
GSM8K MATH Minerva MATHQA ASDiv SVAMP
CodeLlama 12.4 6.0 5.20 14.10 50.5 44.5
CodeLlama-CPT 28.9 11.1 9.80 24.0 61.1 56.4
CodeLlama-Velocitune 28.4 11.7 11.4 25.1 60.9 56.1
Model Math Code
MMLU-STEM SAT Math Avg. HumanEval MBPP Code Avg.
CodeLlama 20.9 18.8 21.6 30.50 43.20 36.80
CodeLlama-CPT 36.0 46.9 34.3 26.20 44.80 35.50
CodeLlama-Velocitune 37.3 56.2 35.9 (+1.6%) 34.10 44.40 39.3 (+3.8%)
Table 1: Performance of CodeLlama-7B, the baseline, and Velocitune on multiple math and code benchmarks.
3.1 Experimental setup
Training corpus
We compile the reasoning dataset based on Proof-Pile-2 (Azerbayev et al., 2023) which consists of math reasoning text in natural language, format language, and code. The dataset included three domains: Arxiv, AlgebraicStack, and OpenWebMath (Paster et al., 2023). Following the common practice to add replay data to prevent catastrophic forgetting, we add two more domains: general code and general language which are composed of the Github subset from SlimPajama (Soboleva et al., 2023) and a blend of Slimpajama except Github and Arxiv. This results in a training set spanning five domains, with 76% of the data in natural language and 24% in code. For system knowledge, we use a dataset SystemStack which is collected from three sources Arxiv, Blogs, and Stackoverflow, concentrating on computer system-related knowledge. The statistics of the two datasets are shown in Table 7.
Training Setup
We trained the models using the standard next token prediction Negative Log-Likelihood (NLL) loss function. Three settings was compared, Velocitune, Dynamic Batch Loading (DBL) (Xia et al., 2023) which update weight using the distance between the evaluation loss and the target loss, and a baseline named CPT which is based on the number of tokens in each domain. For DBL, we also apply the predicted target loss. The total number of tokens processed during training was equivalent for the three methods to completing one full epoch using the training dataset. Detailed hyperparameters for the training process are summarized in Table LABEL:tab:training_details.
Evaluation
To evaluate the math reasoning ability of models, we use math-lm-eval [^1] https://github.com/ZubinGou/math-evaluation-harness from ToRA (Gou et al., 2023) to evaluate accuracy on GSM8K (Cobbe et al., 2021), MATH (Hendrycks et al., 2021), Minerva (Lewkowycz et al., 2022),MMLU-STEM (Hendrycks et al., 2020), ASDiv (Miao et al., 2020), SVAMP (Patel et al., 2021), and MathQA (Amini et al., 2019). For coding ability, we use the evaluation kit from DeepSeek-Coder (Guo et al., 2024) to assess the pass@1 accuracy on HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021).
For evaluating command generation ability, we use two benchmarks CmdGen-NVIDIA and CmdGen-AMD from the CmdGen series, which are built to assess models’ ability to provide proper system command when asked by a related question. The two benchmarks, provided a combination of 1.5K instruction-tuning data and 205 and 192 test questions respectively, evaluate models output from six metrics: Similarity of command(CMD Sim): cosine similarity of CMD embeddings. Similarity of execution output(Output Sim): cosine similarity of output embeddings. Approximate accuracy: CMD Sim or Output Sim is larger than 0.9. The embeddings are generated by all-MiniLM-L6-v2222https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2. Exact Match(EM): If model-generated commands are identical to the target commands. Success Ratio(SR): If the execution of the generated commands incur the same output from the system as the target command. Accuracy(Acc): The union of EM and SR. Example questions and commands are shown in Appendix A.
3.2 Reasoning training results
Velocitune learns math and keeps coding ability.
In Table 1, we list the the nine benchmarks tested for comparing the methods on CodeLlama-7B with Reasoning. Velocitune leads the CPT baseline by 1.9% on seven math benchmarks and 3.8% on code average while baseline’s coding ability dropped by 1.3% from before the continued pre-train. It underscores the reason behind Velocitune’s effectiveness, is balancing the learning velocity in each domain, while in the baseline due to the static domain mixture, the model might learn some domains very well while not saturate for other domains. Velocitune by align the learning velocity across domains, resulting in a more balanced learning progress, thus developing a model better in downstream tasks.
Figure 3: Domain weights dynamic of Velocitune and DBL in training Llama-3 8B on SystemStack.
Table 2: Results of Velocitune, DBL, and basline on Llama3 and Mistral on CmdGen-NVIDIA and CmdGen-AMD benchmarks.
Model CmdGen-NVIDIA
Cmd Sim Output Sim Approx Acc EM SR Acc
Mistral 80.57 58.67 61.95 24.88 19.02 30.73
Mistral-CPT 83.36 65.26 66.34 23.90 21.46 32.20
Mistral-DBL 79.83 65.97 59.02 24.39 19.02 32.20
Mistral-Velocitune 82.85 64.30 68.29 27.32 21.95 36.59 (+4.4%)
Llama-3 86.41 69.09 64.39 41.95 32.68 50.73
Llama-3-CPT 87.53 72.15 69.27 46.34 37.07 57.07
Llama-3-DBL 83.83 64.87 63.90 38.54 27.80 45.85
Llama-3-Velocitune 89.37 75.59 76.59 51.21 39.02 61.95 (+5.3%)
Model CmdGen-AMD
Cmd Sim Output Sim Approx Acc EM SR Acc
Mistral 84.25 59.49 61.54 25.64 15.90 29.23
Mistral-CPT 87.58 65.72 70.77 25.13 18.46 30.77
Mistral-DBL 84.61 67.37 65.13 26.15 17.95 32.82
Mistral-Velocitune 86.84 62.99 69.23 27.18 18.46 33.33 (+2.6%)
Llama-3 88.08 71.22 67.18 41.54 27.69 47.18
Llama-3-CPT 88.69 69.60 70.26 46.15 29.23 51.79
Llama-3-DBL 88.80 68.00 69.23 39.49 27.69 45.64
Llama-3-Velocitune 89.23 72.97 74.36 49.74 31.28 54.87 (+3.1%)
Velocitune dynamically regularises learning velocity.
As shown in Figure 2, the overall weight adjustments can be divided into three distinct stages. During the initial phase of training, the weights for newly introduced domains—OpenWebMath, Arxiv, and AlgebraicStack—rise, while those for the replay domains, GenCode and GenLanguage, decline sharply. This occurs because the LMs are typically less saturated in the new domains, leading to slower learning compared to the replay domains. As a result, Velocitune increase the weights of the underperforming domains to balance their learning. In the second stage, the weights for the replay domains begin to increase while those of the new domains decrease. This shift happens after the models have made substantial progress in the new domains, prompting Velocitune to reallocate focus toward the replay domains. In the final stage, the replay domains’ weights decrease again, while the downslopes for AlgebraicStack and OpenWebMath become more gradual, even showing a slight peak before continuing to decline. This suggests that the model has learned these domains well. Meanwhile, Velocitune increases the weight of Arxiv, where the learning velocity remains slower compared to the others. By adjusting dynamically, Velocitune aligns the learning velocity across domains, ensuring a more balanced learning outcome.
3.3 SystemStack training results
Table 3: Performance comparison of CodeLlama-7B trained using reweighted domain ratios, original domain ratios,Velocitune, and Velocitune without target loss on math and code reasoning benchmarks.
Model Math
GSM8K MATH Minerva MATHQA ASDiv SVAMP
CodeLlama-CPT 28.9 11.1 9.80 24.0 61.1 56.4
CodeLlama-Reweighted 28.7 11.0 12.4 23.6 60.9 56.3
CodeLlama-Velocitune 28.4 11.7 11.4 25.1 60.9 56.1
Model Math Code
MMLU-STEM SAT Math Avg. HumanEval MBPP Code Avg.
CodeLlama-CPT 36.0 46.9 34.3 26.2 44.8 35.5
CodeLlama-Reweighted 37.7 62.5 36.6 30.5 42.8 36.6
CodeLlama-Velocitune 37.3 56.2 35.9 34.1 44.4 39.3
Table 4: Performance comparison of Llama3 models trained reweighted domain ratios, original domain ratios, and Velocitune on the CmdGen benchmarks
Model CmdGen-NVIDIA
Cmd Sim Output Sim Approx Acc EM SR Acc
Llama-3-CPT 88.69 69.60 70.26 46.15 29.23 51.79
Llama-3-Reweighted 87.46 72.09 67.32 46.83 35.61 56.10
Llama-3-Velocitune 89.37 75.59 76.59 51.21 39.02 61.95
Model CmdGen-AMD
Cmd Sim Output Sim Approx Acc EM SR Acc
Llama-3-CPT 88.69 69.60 70.26 46.15 29.23 51.79
Llama-3-Reweighted 88.92 73.16 69.74 47.18 32.31 54.87
Llama-3-Velocitune 89.23 72.97 74.36 49.74 31.28 54.87
Velocitune brings improvements across benchmarks.
On SystemStack, Velocitune improves downstream task performance over DBL and CPT on Llama3-8B and Mistral-7B. Table 2 shows that Velocitune improves the performance in CMD-NVIDIA by 5.3% and 4.4% compared to baseline on Llama3-8B and Mistral-7B, and 3.1% and 2.6% in CMDGen-AMD. It is interesting to find that on Llama3-DBL underperforms baseline in both benchmarks which could attribute to an imbalanced learning brought by updating domain weights based on the distances to the target loss.
Velocitune accelerates weight stablisation
Velocitune accelerates weight stablisation for system domain. We plot the weights moving trajectory on SystemStack comparing Velocitune, DBL, and the baseline in Figure 3. In parallel with Xia et al. (2023), we also observe that the weights stabilize after a few thousand steps. As re-scaling accelerates convergence in machine learning (Juszczak et al., 2002), we also observe that Velocitune on SystemStack reaches stablisation at least 1.5x faster than DBL as DBL does not stablise at the end of training. Observing from the trend, the curves of DBL seem to be a prolonged version of Velocitune. We hypothesis that if the DBL training continues for more steps, it will reach the same final weights similar to Velocitune. For training on Reasoning, we do not observe a stablisation but only the curve becoming towards flat at the end of training.
3.4 Data ordering contribute along with reweight domain weights
To isolate the key contributors behind the effectiveness of Velocitune, we further conduct an ablation study. Specifically, we collect the weights during Velocitune for every evaluation interval, then average over the weights to have the overall sampling ratio of each domain across the training. A comparison of the originial weight and the reweighted weights are shown in Table 9 and 10. Provided with the reweighted data ratio, we train the model on the dataset constantly on the mixture for the same training steps to examine its effect. We conduct the experiment for CodeLlama-7B on Reasoning and Llama3 on SystemStack. The results are shown in Table 3 and Table 4.
Table 5: Performance comparison of CodeLlama-7B with or without target loss.
Model Math
GSM8K MATH Minerva MATHQA ASDiv SVAMP
Velocitune w/o target loss 17.5 8.2 7.4 31.5 43.8 48.8
CPT 28.9 11.1 9.80 24.0 61.1 56.4
Velocitune 28.4 11.7 11.4 25.1 60.9 56.1
Model Math Code
MMLU-STEM SAT Math Avg. HumanEval MBPP Code Avg.
CPT 36.0 46.9 34.3 26.2 44.8 35.5
Velocitune w/o target loss 54.2 30.2 30.2 28.0 42.0 35.0
Velocitune 37.3 56.2 35.9 34.1 44.4 39.3
On the CmdGen benchmarks, the Llama-3 trained with reweighted ratio data outperform Llama3-CPT on both benchmarks, but inferior to Velocitune on NVIDIA and tie on AMD. On math and code evaluation, the CodeLlama trained with reweighted ratio achieves the same average score with Velocitune on the math benchmarks, but still fall behinds Velocitune by 2.7% for coding benchmarks. In general, reweighted data ratios improve downstream task performance compared to the original mixture. However, models still underperform relative to Velocitune, with the sole exception being the Math average, where performance is 0.7% higher than Velocitune. This finding is particularly intriguing, as data mixture is typically assumed to be the primary factor influencing downstream performance. However, this result emphasizes the critical role of dynamic data mixing rather than relying on a static mixture. Previous studies (Hu et al., 2024; AI@Meta, 2024) have highlighted the effects of presenting different data at various stages of model training. Our comparison further supports the potential of dynamic data weighting during pre-training. We encourage further research into the impact of data ordering and its role in optimizing model performance.
3.5 Target loss is critical to Velocitune
To highlight the significance of the target loss, we also explore an alternative method of updating the weights by re-scaling the evaluation losses during training. Specifically, we divide the evaluation loss by the initial loss for each domain, as given by: $\delta^{i}=\frac{\ell^{i}_{T}}{\ell^{i}_{\text{init}}}$. The domain weight dynamic of this is shown in Figure 4 and the corresponding results are presented in Table 5. We see that without the constraints of target loss, scaling solely on the initial loss constantly drives the weight of GenLangauge to nearly 1 and surpress others to 0. Consequently, the model exhibits a performance decrease of 4.1% in the math benchmarks and 0.5% in the code benchmarks compared to the CPT baseline. We hypothesize that this update function implicitly assumes a target loss of zero for each domain, leading to unrealistic expectations for the model to minimize the loss to zero across all domains. Such impractical targets disrupt the training process, ultimately degrading the model’s overall performance.
4 Related Work
GLaM (Du et al., 2022) manually assigns weights to each domain based on the amount of data and the downstream performance of a small model trained on the domain data. Lemma (Azerbayev et al., 2023), an effort to continue pretraining LMs with math-related data from three domains, determines domain weights by iterating through different combinations and selecting those with the lowest perplexity on an evaluation dataset. Doremi (Xie et al., 2024) employs a three-stage process to dynamically adjust domain weights: (1) Train a reference model, (2) Train a proxy model while adjusting the proportion of data based on the proxy model’s loss, and (3) Train the final model using the aggregated data weights, demonstrating superior performance compared to models trained with the original weights. Xia et al. (2023) shows that instead of training a reference model, the target loss value can be estimated using scaling laws (Hoffmann et al., 2022), achieving better performance on smaller LMs with the predicted losses.
5 Conclusion
In this work, we introduced Velocitune, a novel approach for dynamically adjusting domain weights during the training of large language models. Our method addresses the limitations of existing techniques that rely on static or purely loss-based adjustments, which can lead to imbalanced learning across different domains. By scaling to align learning velocity across domains, Velocitune ensures a more balanced learning process, thereby enhancing the model’s ability to generalize across diverse tasks. Our experiments on CodeLlama 7B and Llama3 demonstrate the effectiveness of Velocitune in improving performance on a variety of tasks, including math reasoning, coding, and system diagnosis. Furthermore, our ablation study reveals that the benefits of Velocitune are not solely due to the adjustment of domain weights, but also stem from the synergistic effect of the training dynamics, including the order in which data from different domains is processed. This finding underscores the importance of considering both the weights and the order of learning in the pre-training of large language models.
6 Limitations
The limitations of our study are twofold. Firstly, Velocitune only applies to continual pre-training, and is yet to be examined in pre-training from scratch. Secondly, despite the multi-domain nature of prevelant supervised fine-tuning(SFT) datasets, the possibility of applying Velocitune in SFT has not been explored.
Appendix A Examples of CmdBen
CMDGen-NVIDIA Question: How to compile a CUDA file with a restriction on register count per thread to 32 and enable verbose output? Target Command: nvcc -v -maxrregcount=32 sample.cu
CMDGen-AMD Question: How to use ’rocm-smi’ to show the range of memory clock frequencies (mclk) that can be set? Target Command: rocm-smi –showmclkrange
Appendix B Scaling law prediction error
Table 6: The average error over domain for predicting evaluation loss of models using full training data. We used evaluation loss from checkpoints saved till using half the training data to fit the laws.
Reasoning SystemStack
Error 2.4e-3 1.84e-3
Appendix C Training datasets statistics
Table 7: Statistics for Proof-Pile-2Plus and SystemStack. Token counts were determined using the CodeLlama and Llama3 tokenizers, respectively.
Reasoning
Domain # Tokens (Billion)
Arxiv 28.70
AlgebraicStack 10.47
OpenWebMath 14.02
GenCode 3.97
GenLangauge 2.92
SystemStack
Domain # Tokens (Billion)
Arxiv 5.37
Blogs 3.21
StackOverflow 7.64
Appendix D Training Details
We ran the experiments on 64 Nvidia H100s.
Table 8: Training hyper-parameters and throughput.
SystemStack Reasoning
Training tokens $16$B $63$B
Learning rate $51e-$ 5 $55e-$ 5
LR warmup ratio $0.005$ $0.005$
Batch size (tokens) $1$M $4$M
Evaluation interval $m$ (steps) $150$ $150$
Steps $15,482$ $13,807$
# GPUs $64$ $64$
Sample ratio for proxy model 58% 51%
Adam $\beta_{2}$ 0.95 0.99
Appendix E Reweighted Data Ratio after Velocitune
The reweighted data ratio is shown in Table 3 and Table 9.
Table 9: Original and Velocitune reweighted ratio on SystemStack for Llama3
Domain Blogs Stackoverflow Arxiv
Original 0.187 0.503 0.310
Reweighted 0.197 0.470 0.333
Table 10: Original and Velocitune reweighted ratio on Reasoning for CodeLlama
Domain AlgebraicStack Arxiv OpenWebMath GenCode GenLanguage
Original 0.189 0.500 0.259 0.029 0.020
Reweighted 0.185 0.463 0.225 0.086 0.040
Appendix F Domain weight dynamic in Velocitune without target loss
Figure 4: Domain weights dynamic of Velocitune without target loss in training CodeLlama 7B on Math&Code.